# On the Structure and Semantics of Identifier Names Containing Closed Syntactic Category Words

### v1

@@ Abstract
Identifier names are crucial components of code, serving as primary clues for developers to understand program behavior. This paper investigates the linguistic structure of identifier names by extending the concept of grammar patterns; representations of the part-of-speech (PoS) sequences that underlie identifier phrases. The specific focus is on closed syntactic categories (e.g., prepositions, conjunctions, determiners), which are rarely studied in software engineering despite their central role in general natural language. The Closed Category Identifier Dataset (CCID) is presented, a new manually annotated dataset of 1,275 identifiers drawn from 30 open-source systems. The relationship between closed-category grammar patterns and program behavior is analyzed using grounded theory coding, statistical, and pattern analysis. The results reveal recurring structures that developers use to express control flow, data transformation, temporal reasoning, and behavioral roles through naming. This study contributes an empirical foundation for understanding how developers adapt linguistic resources to encode behavior in source code. By analyzing closed-category terms and their associated grammar patterns, the work highlights a previously underexplored dimension of identifier semantics and identifies promising directions for future research in naming support, comprehension, and education.
@@ 1 Introduction
At a high level, PoS can be split into two Syntactic Categories: open and closed. Most identifier naming research has focused on open-category, which includes nouns and verbs. The set of open category terms changes and increases (in terms of new words) with time as new domains emerge and change. In contrast, closed-category (e.g., prepositions, conjunctions, determiners) are drawn from a fixed set and serve functional roles in language; this set of terms rarely sees new words introduced over time. These terms have received little attention in the software literature, despite their importance in human languages. Identifying closed-category terms in code is also non-trivial: for example, the word and may represent a conjunction or a logical operator, depending on context—making PoS tagging a prerequisite for meaningful analysis.
@@
  • A mixed-methods analysis combining grounded theory coding with statistical evaluation to characterize the semantics of closed-category grammar patterns and their constituent terms.
@@
Our findings have implications for both human and automated naming support. Grammar patterns offer a structured way to analyze naming behavior, surface potential inconsistencies, and guide naming suggestions. For AI-based tools, they offer scaffolding to align generated names with human conventions. For developers and educators, they reveal naming idioms that can support clearer communication and pedagogy. In this study, we address the following research questions:
RQ1: What behavioral roles do closed-category terms play in source code identifiers? To address this question, we conduct a grounded theory study on a manually annotated dataset of identifiers containing closed-category terms: prepositions, conjunctions, determiners, and digits. Through iterative coding and memoing, we develop axial and selective codes that describe the behavioral functions these terms convey in source code, such as data flow, condition handling, or execution sequencing. This process allows us to uncover not only common grammar patterns but also the communicative intent behind developers’ use of closed-category terms. Our goal is to characterize the nuanced and purposeful ways in which these terms encode program behavior.
RQ2: How do closed-category terms correlate with structural, programming language, and domain-specific contexts in software? To answer this question, we quantitatively analyze the distribution of closed-category terms across multiple dimensions: source-code-local structure (e.g., function names, parameters, class names), programming languages (e.g., Java, C++, C), and system domains (e.g., libraries, frameworks, domain-specific applications). We use statistical tests to examine whether these terms appear disproportionately in certain contexts. These correlations help us determine whether developers systematically leverage closed-category terms to express behavior in ways that are shaped by structural conventions, linguistic norms, or domain constraints.
This paper is organized as follows. Section 2 provides our reasoning on why it is important to study this topic. Section 3 gives background on grammar patterns in the context of identifier names. Section 4 provides a detailed explanation of our methods for undertaking the investigation. Our Evaluations are presented in Sections 5 and 6. Related work on identifier names is in Section 7. Discussion of the results is in section 8, followed by Threats to Validity in Section 9. Conclusions are in Section 10 and Data Availability in Section 11.
@@ 2 Why Study Closed-Category Naming Patterns?
Despite their potential significance, these terms have received almost no attention in prior software development naming research, which has focused primarily on open-category words (e.g., nouns, verbs). As a result, we lack foundational knowledge about when and how closed-category terms are used in code—and what they contribute to program comprehension.
Understanding these naming patterns has clear implications: it can inform naming tools, guide educational resources, improve automated name generation, and help researchers characterize naming conventions more precisely. Closed-category terms may be uncommon, but we argue, in this paper, that their usage is not accidental; they significantly contribute to the meaning of identifier names, making it important to study them.
@@ 3 Definitions & Grammar Pattern Generation
In this work, we analyze identifier names through the lens of grammar patterns, which are sequences of part-of-speech (PoS) tags assigned to the terms within an identifier. For example, the identifier GetUserToken is split into the terms Get, User, and Token, which are tagged as Verb Noun-adjunct Noun. This sequence, V NM N, represents the identifier’s grammar pattern. Crucially, this pattern generalizes across many identifiers: RunUserQuery and WriteAccessToken share the same structure, even though they use different terms. Grammar patterns thus allow us to relate identifiers by their syntactic form.
We focus specifically on closed-category grammar patterns, which are patterns that contain at least one closed-class part of speech: a preposition, determiner, conjunction, or numeral (digit). These categories are finite and rarely accept new terms, in contrast to open-class categories like nouns and verbs, which grow over time as new domains introduce new concepts. Despite their rarity in code, closed-category terms often signal behavioral relationships such as event triggers, quantification, or conditional logic, making them important to study.
@@ 3.1 Phrasal Structures and Interpretation
Repo Link Name Primary Language Commit Hash Data of most recent commit C LOC C++ LOC Java LOC Total LOC
https://github.com/liuliu/ccv ccv C 3eb1664f88e04515bf6861663c4f96d6279a47f5 2023-07-19 279186 1908 0 281094
https://github.com/ropensci/git2r git2r C 9c42d41c14cf83735db37ae662e2bce42c500aaa 2023-05-01 99956 0 0 99956
https://github.com/Juniper/libxo libxo C 290d88625c683bca60e05e68dd2003c290f50f99 2023-02-08 9361 0 0 9361
https://github.com/mgba-emu/mgba mgba C c8cfaefcc880e668b0db8ba5f8b49978f0d108fd 2023-07-18 302865 25309 0 328174
https://github.com/naemon/naemon-core naemon-core C 7558c98937ffb15c040daf158f92827b1f338ba6 2023-07-07 40991 0 0 40991
https://github.com/openvswitch/ovs ovs C f5188ff2147517612b410ed607e3843cdf4b51a6 2023-07-19 268376 0 0 268376
https://github.com/igraph/rigraph rigraph C cae1f110975e38353a8fbbe16d4c7a008fbd9bea 2023-07-19 273973 32737 0 306710
https://github.com/toggl-open-source/toggldesktop toggldesktop C 013038304ee8e9154db4a45fb642716956cf153b 2023-06-22 582087 269105 0 851192
https://github.com/irungentoo/toxcore toxcore C bf69b54f64003d160d759068f4816b2d9b2e1e21 2018-10-03 27443 0 0 27443
https://github.com/weechat/weechat weechat C 13241cdca5b39415d2a0f9e1b55a79153d14ba07 2023-07-20 197608 31175 0 228783
https://github.com/wireshark/wireshark wireshark C 13f21ebd59317f59b91ce7a48a234ef48989ccf0 2023-07-20 4171790 102848 0 4274638
https://github.com/BVLC/caffe caffe C++ 9b891540183ddc834a02b2bd81b31afae71b2153 2020-02-13 0 42856 0 42856
https://github.com/vgvassilev/cling cling C++ 5e3310ac045d5763cd9852359d1003f3850c2a0c 2023-07-18 57 28342 0 28399
https://github.com/ipkn/crow crow C++ 2b43d3cd6a9a9cdbc99dfef9b86ff3f3027f3d1f 2022-09-20 0 1434 0 1434
https://github.com/fakeNetflix/facebook-repo-ds2 ds2 C++ eaf83c59b8fff11cc4c218da7b78a983754727ec 2019-07-17 367 27011 0 27378
https://github.com/freeminer/freeminer freeminer C++ 86caf1cfbdf3d06f53981adaedc6f1eba7448486 2023-04-22 15090 130828 1077 146995
https://github.com/meta-toolkit/meta meta C++ c7019401185cdfa15e1193aad821894c35a83e3f 2017-08-19 132 25451 0 25583
https://github.com/panda3d/panda3d panda3d C++ 07f9f9d89722a82234e0d111e61b7ff98e88e2db 2023-06-13 44671 416212 175 461058
https://github.com/facebook/proxygen proxygen C++ 5370e58851d6b363f1edbc542a0daa369ebbf303 2023-07-20 2676 70161 0 72837
https://github.com/s3fs-fuse/s3fs-fuse s3fs-fuse C++ 5371cd1468c84423729c334ac997f9621b797e9f 2023-07-19 197 19582 0 19779
https://github.com/cglib/cglib cglib Java 5c947539c1218dc4ff1a14a7c36b4856b0ee6397 2022-02-08 0 0 15187 15187
https://github.com/deeplearning4j/deeplearning4j deeplearning4j Java 854c3b264d6d392ecd1b09c386d919e3c74cb425 2023-06-21 0 224997 696717 921714
https://github.com/apache/drill drill Java f558815feb6560ebbc3d1fcf502acee8fea1e555 2023-06-21 538 34591 626295 661424
https://github.com/google/guava guava Java 551f92130cc0442870569b89820e0b432a8ff8f7 2023-07-18 0 0 356651 356651
https://github.com/immutables/immutables immutables Java 28ba97e9f3075902e279c1cde3e16be01e365afb 2023-06-16 0 0 69505 69505
https://github.com/dropwizard/metrics metrics Java d6cf6a4e8f6e65a0b6107be6d8d73e00f0d95efa 2023-07-20 0 0 31317 31317
https://github.com/igniterealtime/Openfire Openfire Java 25ed0d66f9c56f423f9408920ce7d8734dcaf84d 2023-07-20 120 0 122186 122306
https://github.com/HubSpot/Singularity Singularity Java 15688f486fc9286878eff40b51789c88bd6899d5 2022-11-18 0 0 122183 122183
https://github.com/igniterealtime/Smack Smack Java 19b20fefeca69cccc3637e01ec2f151a054a3351 2023-04-26 0 0 125547 125547
https://github.com/igniterealtime/Spark Spark Java ca028f3a818b94c00f2f600933109087d84fb0b9 2023-05-11 9 0 91886 91895
@@ 4 Methodology
For our study, identifiers are collected from and analyzed in the following contexts: class names, function names, parameter names, attribute names (i.e., data members), and declaration-statement names. A declaration-statement name is a name belonging to a local (to a function) or global variable. We use this terminology because it is consistent with srcML’s terminology collard:2016 for these variables, and we used srcML to collect identifiers. Therefore, to study closed-category grammar patterns, we group identifiers based on these five categories. The purpose of doing this is to study closed-category grammar pattern frequencies based on their high-level semantic role (e.g., class names have a different role than function names). We collected these identifiers from 30 open source systems, which can be found in Table 3. These systems belonged to a curated dataset of engineered software projects, synthesized by Reaper Munaiah2017 , which is a tool that measures how well different projects follow software engineering practices such as documentation and continuous integration.
The set of systems have an average and median of 335,358 and 111,069 LOC, respectively. 11 of the systems are primarily C systems, 9 are primarily C++, and 10 are primarily Java. We chose systems that have tests and use continuous integration (CI) under the idea that these represented systems that have at least some basic process for ensuring quality; Reaper is able to automatically determine which systems have both CI and tests. Our primary concern for selecting systems is that they represent different programming languages, follow basic quality procedures, and are large enough for us to collect enough identifiers. Given this, our choice of systems is designed to ensure that the grammar patterns in this study are applied across at least the languages under study.
@@ 4.1 Detecting and Sampling Identifiers with Closed-Category Terms
Sampling identifiers that contain closed-category terms is challenging for two reasons: (1) They are relatively uncommon in production code, and (2) many such terms are ambiguous without context, making automatic tagging difficult. To address this, we implemented a two-phase sampling strategy: (1) filtering identifiers that potentially contain closed-category terms into candidate sets, and (2) manually verifying and annotating a statistically representative sample.
@@
  • Digits (D): We selected identifiers containing at least one numeric digit, using Python’s isDigit() functionality. Digits are easier to detect automatically and unambiguous in token form. However, there are cases where a digit will be annotated as part of another category. For example, str2int uses the digit 2 as a preposition (to).
@@
  • 4,630 identifiers candidate digits
We use the term candidate because these filters do not account for context or usage, and thus include false positives. Still, they serve as an upper bound for each category’s prevalence in the corpus. Based on this, we estimate the proportion of identifiers containing each term type as follows:
@@
  • D: 355 candidate digit identifiers
@@
Some contexts, such as parameters and especially class names, were underrepresented due to the natural scarcity of closed-category terms in those positions. To increase representation from underrepresented contexts, we attempted to oversample relevant subgroups. However, even with oversampling, the absolute number of qualifying identifiers (e.g., a parameter containing a conjunction) remained low. As our sampling was driven by the presence/population of closed-category terms in general, rather than population within our program contexts under study, we opted not to artificially balance the dataset further.
@@
  • 363 candidate digit identifiers
@@
Following manual verification and annotation (described in Section 4.2), we retained only those identifiers that were confirmed to contain closed-category terms. Table 5 summarizes both the sampled totals and the verified counts. The final dataset consists of 1,001 identifiers confirmed to contain at least one closed-category term. These identifiers comprise the CCID. Table 6 shows the CCID, but broken down by program context instead of closed-category type.
Annotation Notes
Some tools exist for part-of-speech tagging of source code identifiers (e.g., the ensemble tagger from newman_ensemble ), but these are slow at scale and are trained on datasets that underrepresent closed-category terms. For example, in our prior dataset Newman_GP , used to train the aforementioned tagging approach newman_ensemble ; conjunctions, determiners, and prepositions made up only 0.2%, 0.4%, and 2.6% of tags, respectively. Thus, we found manual annotation necessary to ensure sufficient coverage and correctness for our study.
As we are primarily concerned with production code, and prior work shows that test name grammar patterns differ from production names peruma_test_name_gp , we did not collect any identifier containing the word ‘test’, or that appeared in a clearly marked test file or directory. In addition, note that Table 4 counts tags at the word level (e.g., CJ CJ N counts two CJ tags), whereas Table 5 counts tags at the identifier level (e.g., one identifier with multiple CJ tags counts as one). This explains occasional mismatches between sampled and actual tag distributions.

### v4

@@ Abstract
Identifier names are crucial components of code, serving as primary clues for developers to understand program behavior. This paper investigates the linguistic structure of identifier names by extending the concept of grammar patterns, which represent the part-of-speech (PoS) sequences underlying identifier phrases. The specific focus is on closed syntactic categories (e.g., prepositions, conjunctions, determiners), which are rarely studied in software engineering despite their central role in general natural language. To study these categories, the Closed Category Identifier Dataset (CCID), a new manually annotated dataset of 1,275 identifiers drawn from 30 open-source systems, is constructed and presented. The relationship between closed-category grammar patterns and program behavior is then analyzed using grounded-theory-inspired coding, statistical, and pattern analysis. The results reveal recurring structures that developers use to express concepts such as control flow, data transformation, temporal reasoning, and other behavioral roles through naming. This work contributes an empirical foundation for understanding how linguistic resources encode behavior in identifier names and supports new directions for research in naming, program comprehension, and education.
@@ 1 Introduction
At a high level, PoS can be split into two Syntactic Categories: open and closed. Most identifier naming research has focused on open-category, which includes nouns and verbs. The set of open category terms evolves and expands (in terms of new words) over time as new domains emerge and evolve. In contrast, closed-category (e.g., prepositions, conjunctions, determiners) are drawn from a fixed set and serve functional roles in language; this set of terms rarely sees new words introduced over time. These terms have received little attention in the software literature, despite their importance in human languages. Identifying closed-category terms in code is also nontrivial: for example, the word and may represent a conjunction or a logical operator, depending on context, making PoS tagging a prerequisite for meaningful analysis.
@@
  • A mixed-methods analysis combining grounded-theory-style coding with statistical evaluation to characterize the semantics of closed-category grammar patterns and their constituent terms.
@@
Our findings have implications for both human and automated naming support. Grammar patterns provide a structured approach to analyzing naming behavior, identifying potential inconsistencies, and providing naming suggestions. For AI-based tools, they offer scaffolding to align generated names with human conventions. For developers and educators, they reveal naming idioms that can support clearer communication and pedagogy. In this study, we address the following research questions:
RQ1: What behavioral roles do closed-category terms play in source code identifiers? To address this question, we conducted a grounded-theory-inspired study on a manually annotated dataset of identifiers containing closed-category terms: prepositions, conjunctions, determiners, and numerals. Through open coding and memoing, we develop axial and selective codes that describe the behavioral functions these terms convey in source code, such as data flow, condition handling, or execution sequencing. This process allows us to uncover not only common grammar patterns but also the communicative intent behind developers’ use of closed-category terms. Our goal is to characterize the nuanced and purposeful ways in which these terms encode program behavior and convey information.
RQ2: How do closed-category terms correlate with structural, programming language, and domain-specific contexts in software? To answer this question, we quantitatively analyze the distribution of closed-category terms across multiple dimensions: source-code-local structure (e.g., function names, parameters, class names), programming languages (e.g., Java, C++, C), and system domains (e.g., libraries, frameworks, domain-specific applications). We use statistical tests to examine whether these terms appear disproportionately in specific contexts. These correlations help us determine whether developers systematically leverage closed-category terms to express behavior in ways that are shaped by structural conventions, linguistic norms, or domain constraints.
This paper is organized as follows. Section 2 provides our reasoning on why it is essential to study this topic. Section 3 gives background on grammar patterns in the context of identifier names. Section 4 provides a detailed explanation of the methods used for conducting the investigation. Our Evaluations are presented in Sections 5 and 6. Related work on identifier names is in Section 7. Discussion of the results is in section 8, followed by Threats to Validity in Section 9. Conclusions are in Section 10 and Data Availability in Section 11.5.
@@ 2 Why Study Closed-Category Naming Patterns?
Despite their potential significance, these terms have received almost no attention in prior software development naming research, which has focused primarily on open-category words (e.g., nouns, verbs). As a result, we lack foundational knowledge about when and how closed-category terms are used in code and what they contribute to program comprehension.
Understanding these naming patterns has clear implications: it can inform naming tools, guide educational resources, improve automated name generation, and help researchers characterize naming conventions more precisely. Closed-category terms may be uncommon, but we argue, in this paper, that their usage is not accidental; they significantly contribute to the meaning of identifier names, making it important to study them. More examples of identifiers containing closed-category terms can be found in Table 1.
@@ 3 Definitions & Grammar Pattern Generation
In this work, we analyze identifier names through the lens of grammar patterns, which are sequences of part-of-speech (PoS) tags assigned to the terms within an identifier. For example, the identifier GetUserToken is split into the terms Get, User, and Token, which are tagged as Verb Noun-adjunct Noun. This sequence, V NM N, represents the identifier’s grammar pattern. Crucially, this pattern generalizes across many identifiers: RunUserQuery and WriteAccessToken share the same structure, despite using different terms. Grammar patterns thus allow us to relate identifiers by their syntactic form.
We focus specifically on closed-category grammar patterns, which are patterns that contain at least one closed-class part of speech: a preposition, determiner, conjunction, or numeral. These categories are finite and rarely accept new terms, in contrast to open-class categories like nouns and verbs, which grow over time as new domains introduce new concepts. Despite their rarity in code, closed-category terms often signal behavioral relationships such as event triggers, quantification, or conditional logic, making them critical to study.
@@ 3.1 Phrasal Structures and Interpretation
Repo Link Name Primary Language Data of most recent commit C LOC C++ LOC Java LOC Total LOC
https://github.com/liuliu/ccv ccv C 2023-07-19 279186 1908 0 281094
https://github.com/ropensci/git2r git2r C 2023-05-01 99956 0 0 99956
https://github.com/Juniper/libxo libxo C 2023-02-08 9361 0 0 9361
https://github.com/mgba-emu/mgba mgba C 2023-07-18 302865 25309 0 328174
https://github.com/naemon/naemon-core naemon-core C 2023-07-07 40991 0 0 40991
https://github.com/openvswitch/ovs ovs C 2023-07-19 268376 0 0 268376
https://github.com/igraph/rigraph rigraph C 2023-07-19 273973 32737 0 306710
https://github.com/toggl-open-source/toggldesktop toggldesktop C 2023-06-22 582087 269105 0 851192
https://github.com/irungentoo/toxcore toxcore C 2018-10-03 27443 0 0 27443
https://github.com/weechat/weechat weechat C 2023-07-20 197608 31175 0 228783
https://github.com/wireshark/wireshark wireshark C 2023-07-20 4171790 102848 0 4274638
https://github.com/BVLC/caffe caffe C++ 2020-02-13 0 42856 0 42856
https://github.com/vgvassilev/cling cling C++ 2023-07-18 57 28342 0 28399
https://github.com/ipkn/crow crow C++ 2022-09-20 0 1434 0 1434
https://github.com/fakeNetflix/facebook-repo-ds2 ds2 C++ 2019-07-17 367 27011 0 27378
https://github.com/freeminer/freeminer freeminer C++ 2023-04-22 15090 130828 1077 146995
https://github.com/meta-toolkit/meta meta C++ 2017-08-19 132 25451 0 25583
https://github.com/panda3d/panda3d panda3d C++ 2023-06-13 44671 416212 175 461058
https://github.com/facebook/proxygen proxygen C++ 2023-07-20 2676 70161 0 72837
https://github.com/s3fs-fuse/s3fs-fuse s3fs-fuse C++ 2023-07-19 197 19582 0 19779
https://github.com/cglib/cglib cglib Java 2022-02-08 0 0 15187 15187
https://github.com/deeplearning4j/deeplearning4j deeplearning4j Java 2023-06-21 0 224997 696717 921714
https://github.com/apache/drill drill Java 2023-06-21 538 34591 626295 661424
https://github.com/google/guava guava Java 2023-07-18 0 0 356651 356651
https://github.com/immutables/immutables immutables Java 2023-06-16 0 0 69505 69505
https://github.com/dropwizard/metrics metrics Java 2023-07-20 0 0 31317 31317
https://github.com/igniterealtime/Openfire Openfire Java 2023-07-20 120 0 122186 122306
https://github.com/HubSpot/Singularity Singularity Java 2022-11-18 0 0 122183 122183
https://github.com/igniterealtime/Smack Smack Java 2023-04-26 0 0 125547 125547
https://github.com/igniterealtime/Spark Spark Java 2023-05-11 9 0 91886 91895
@@ 4 Methodology
For our study, identifiers are collected from and analyzed in the following contexts: class names, function names, parameter names, attribute names (i.e., data members), and declaration-statement names. A declaration-statement name is a name belonging to a local (to a function) or global variable. We use this terminology because it is consistent with srcML’s terminology collard:2016 for these variables, and we used srcML to collect identifiers. Therefore, to study closed-category grammar patterns, we group identifiers based on these five categories. The purpose of this categorization is to examine the closed-category grammar patterns based on their high-level semantic roles (e.g., class names have a different role than function names). We collected these identifiers from 30 open-source systems, which are listed in Table 3. These systems belonged to a curated dataset of engineered software projects, synthesized by Reaper Munaiah2017 , a tool that measures how well different projects adhere to software engineering practices, such as documentation and continuous integration.
The set of systems has an average and median of 335,358 and 111,069 LOC, respectively. 11 of the systems are primarily C systems, 9 are mainly C++, and 10 are primarily Java. We chose systems that have tests and use continuous integration (CI) under the idea that these represent systems with at least some basic process for ensuring quality; Reaper is able to automatically determine which systems have both CI and tests. Our primary concern in selecting systems is that they represent different programming languages, follow basic quality procedures, and are large enough for us to collect a sufficient number of identifiers. Given this, our choice of systems is designed to ensure that the grammar patterns in this study are applied across at least the languages under study.
@@ 4.1 Detecting and Sampling Identifiers with Closed-Category Terms
Sampling identifiers that contain closed-category terms is challenging for two reasons: (1) They are relatively uncommon in production code, and (2) many such terms are ambiguous without context, making automatic tagging difficult. Table 4 shows the distribution of closed-category PoS tags present in a data set we constructed in prior work Newman_GP versus the CCID, which was constructed explicitly to increase the population of closed-category PoS in the set. To address this, we implemented a two-phase sampling strategy: (1) filtering identifiers that potentially contain closed-category terms into candidate sets, and (2) manually verifying and annotating a statistically representative sample.
@@
  • Numerals (D): We selected identifiers containing at least one digit, using Python’s isdigit() functionality. Numerals are easier to detect automatically and are unambiguous in token form. However, there are cases where a numeral will be annotated as part of another category. For example, str2int uses the numeral 2 as a preposition (to).
@@
  • 4,630 identifiers candidate numerals
We use the term candidate because these filters do not account for context or usage, and thus include false positives. Still, they serve as an upper bound for the prevalence of each category in the corpus. Based on this, we estimate the proportion of identifiers containing each term type as follows:
@@
  • D: 355 candidate numeral identifiers
@@
In some contexts, such as parameters and especially class names, terms were underrepresented due to the natural scarcity of closed-category terms in those positions. To increase representation from underrepresented contexts, we attempted to oversample relevant subgroups. However, even with oversampling, the absolute number of qualifying identifiers (e.g., a parameter containing a conjunction) remained low. As our sampling was driven by the presence/population of closed-category terms in general, rather than population within our program contexts under study, we opted not to artificially balance the dataset further.
@@
  • 363 candidate numeral identifiers
@@
Following manual verification and annotation (described in Section 4.2), we retained only those identifiers that were confirmed to contain closed-category terms. Table 5 summarizes both the sampled totals and the verified counts. The final dataset consists of 1,001 identifiers confirmed to include at least one closed-category term. These identifiers comprise the CCID. Table 6 shows the CCID, but broken down by program context instead of closed-category type.
Annotation Notes. Some tools exist for part-of-speech tagging of source code identifiers (e.g., the ensemble tagger from newman_ensemble ), but these are slow at scale and are trained on datasets that underrepresent closed-category terms. For example, in our prior dataset Newman_GP , used to train the aforementioned tagging approach newman_ensemble ; conjunctions, determiners, and prepositions made up only 0.2%, 0.4%, and 2.6% of tags, respectively. Thus, we found manual annotation necessary to ensure sufficient coverage and correctness for our study.
As we are primarily concerned with production code, and prior work shows that test name grammar patterns differ from production names peruma_test_name_gp , we did not collect any identifiers containing the word ‘test’, or that appeared in a clearly marked test file or directory. In addition, note that Table 4 counts tags at the word level (e.g., CJ CJ N counts two CJ tags), whereas Table 5 counts tags at the identifier level (e.g., one identifier with multiple CJ tags counts as one). This explains occasional mismatches between sampled and actual tag distributions.